Explain your understanding of the phrase base 10number system (4marks)


1
Expert's answer
2020-10-20T17:49:55-0400

Base 10 number system is the numbering system most commonly used around the world. In base 10, each digit in a position has 10 possibilities ranging from 0 to 9 (i.e 0, 1, 2, 3, 4, 5, 6, 7, 8, and 9).


Base 10 numbering system is also known as the decimal(denary) numbering system because the value of each digit is determined by its position in relation to the decimal point.


The number nearest the decimal point to the left has the power of one or units. The next to the left has the power of ten, the next the power of hundred, the next thousand, and so on it goes.

Comparably, the number nearest the decimal point to the right has the power of a tenth, the next has the power of a hundredth, the next a thousandth, and so on.


This characteristic of Base 10 digits is called its place value.


For instance, in the number 5,437,967.134;

"5\ has\ a \ place\ value\ of\ 5\, million\,\,\,\ (5,000,000)\\

4\ has\ a \ place\ value \ of\ 4 \ hundred \ thousand\,\,\,\ (400,000)\\

3\ has\ a \ place\ value\ of \ 3\ ten-thousand\,\,\,\ (30,000)\\

7 \ has \ a \ place\ value \ of\ 7\ thousand\,\,\,\ (7,000)\\

9 \ has\ a\ place \ value\ of \ 9\ hundred\,\,\,\ (900)\\

6\ has \ a \ place\ value \ of\ 6\ ten\,\,\,\ (60)\\

7\ has\ a \ place\ of\ 7\\

1\ has\ a \ place\ of\ 1 \ one-tenth\,\,\,\ (\frac{1}{10})\\

3\ has \ a \ place\ value \ of \ 3 \ one-hundredth\,\,\,\ (\frac{3}{100})\\

4\ has\ a \ place\ value\ of \ 4\ one-thousandth\,\,\,\ (\frac{4}{1000})\\"